OpenStudy (anonymous):

Write the exponential function y = 82(1.18)^t in the form y= ae^kt

OpenStudy (aum):

\[\large x = e^{\ln(x)} \\ \large 1.18 = e^{\ln(1.18)} = e^{0.1655} \\ \large y = 82(1.18)^t = 82(e^{0.1655})^t = 82 e^{0.1655t} \]

OpenStudy (aum):

Any number can be written as e^( ln(that number) ) 5 = e^(ln(5))

OpenStudy (anonymous):

oh thank you

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

how would you get the annual percent growth rate and the continous